OpenStudy (anonymous):

If x > 0, then (4^x)(8^x) = ?

OpenStudy (anonymous):

If x > 0 then, \[(4^{x}) (8^{x})\] =

OpenStudy (gorv):

\[4=2^2\] \[8=2^3\] \[(2^2)^x*(2^3)^x\]

OpenStudy (gorv):

(2^2x)*(2^3x)=2^(2x+3x)=2^(5x)

OpenStudy (gorv):

@SlightlyConfused

OpenStudy (anonymous):

thanks! still a bit confused though but thanks

OpenStudy (gorv):

tell me buddy @SlightlyConfused

OpenStudy (anonymous):

I'm confused on this part; (2^2x)*(2^3x)=2^(2x+3x)=2^(5x)

OpenStudy (anonymous):

I got the first part, but how did you go from there to that long equation?

OpenStudy (anonymous):

never mind i got it! you multiplied the 2 and the x!!!!!!!!!!!

OpenStudy (gorv):

if (2^a)^b then powers will be multiplied (2^a)^b=2^(a*b)

OpenStudy (gorv):

now if base is same then power will be added we converted it to a common base 2 so thsat we can add up the powers

OpenStudy (gorv):

@SlightlyConfused ?????

OpenStudy (anonymous):

okay i got it! thanks so much!!!!!!!!!!

OpenStudy (gorv):

alwys welcome :)